# More for Keys, Less for Values: Adaptive KV Cache Quantization

### v1

@@ Appendix A Quantization and Dequantization Error Bounds
Spectral Norm Bound
@@
Frobenius Norm Bound
@@
Spectral-Norm Perspective
@@
Frobenius-Norm (MSE) Perspective
@@
Appendix B Quantization Error Across Layers and Heads
Figure 6 displays the MSE of the quantized key and value caches in the Llama 3.1 8B model, evaluated layer by layer across 32 layers and differentiated by attention head. Results are presented for three datasets: C4, MMLU, and GSM8k. Each subfigure illustrates the 2-bit and 4-bit quantization configurations for the key and value caches. The error patterns vary across layers, suggesting that certain layers or heads may be more sensitive to aggressive quantization; however, the overall pattern indicates that assigning 4-bit to the key cache is comparable to assigning 2-bit to the value cache.
(a) C4 dataset
(b) MMLU dataset
(c) GSM8k dataset
Appendix C Extended Visualization of Singular Value Distributions
@@
This figure underscores two main observations. First, key matrices systematically exhibit higher leading singular values than value matrices. In particular, the first singular value (the spectral norm) for key matrices is consistently larger, confirming that the key matrices possess a higher spectral norm. Second, this behavior (significant gap) remains evident across the entire spectrum of singular values, implying that the sum of the squares of these values (i.e., the Frobenius norm) is also larger for key matrices. Consequently, the stronger representational capacity of key matrices, as reflected by both their spectral and Frobenius norms, renders them more sensitive to quantization errors.
Appendix D Extended Quantization Error Results
Tables 3, 4, and 5 present extended quantization error results across a wider range of models and settings for the MMLU, C4, and GSM8K datasets, respectively. These tables offer a detailed breakdown of the error metrics under various bit-width configurations, complementing the analysis presented in the main text and in Table 1.

### v2

@@ C.1 Preliminaries and Notation
C.2 Spectral Norm Bound
@@
C.3 Frobenius Norm Bound
@@ C.4 Quantizing KV Cache
C.5 Spectral-Norm Perspective
@@
C.6 Frobenius-Norm (MSE) Perspective
@@
Appendix D Supplementary Key-Value Weight Norm Analysis
Here we present additional evidence supporting the Key-Value Norm Disparity theorem established in Section 3.1. While Fig. 2 demonstrates this observation for the Llama model family, the same pattern consistently emerges across Mistral family.
Figure 4: Frobenius norm plot for Mistral Family. The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude. The spectral norms are higher for the key weights than for the value weights across layers.
Appendix E Extended Visualization of Singular Value Distributions
@@
Figure 7: Spectral norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the layer index. The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer. Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
Appendix F Extended Quantization Error Results
@@
Figure 8: Mean squared error (MSE) of KV quantization for Llama 3.3-70B on C4. We use quantization bit-widths ranging from 2 to 8. The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale. A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit. Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
(a) C4 dataset
(b) MMLU dataset
(c) GSM8k dataset
(a) C4 dataset
(b) MMLU dataset
(c) GSM8k dataset
Appendix G Reproducibility and Resources
Inference was performed using the Hugging Face Transformers [53] and Accelerate [54] libraries, together with FlashAttention [55, 56], to optimize token-generation throughput. We integrated both Quanto and HQQ into the Language Model Evaluation Harness [57] to enable systematic and reproducible evaluation of model performance under varying quantization schemes. All runs were executed on GPU clusters equipped with NVIDIA A100 (80 GB) using the Quanto backend and NVIDIA H200 (141 GB) using the HQQ backend.
Appendix H Extended Evaluation Results
Extended evaluation results for KV cache quantization under the HQQ backend are presented in Tables 910 (GSM8K flexible and strict accuracies), Tables 8 (EQ-Parseable parseable-pass rates and exact-match scores), and Table 7 (CoQA F1 scores).
Table 7: Full CoQA Accuracies F1.
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0.129657 0.147672 0.127875 0.222218 0.127204 0.358652 0.272695 0.231077
k1v2 0.226803 0.236351 0.222816 0.360911 0.206763 0.384316 0.329965 0.327598
k1v4 0.226441 0.236751 0.229168 0.375805 0.241965 0.385504 0.323921 0.34914
k1v6 0.214566 0.242716 0.229515 0.378638 0.220753 0.378799 0.34787 0.351566
k1v8 0.219723 0.240167 0.221966 0.364095 0.241671 0.382309 0.341797 0.3545
k2v1 0.252746 0.189839 0.576534 0.463689 0.565141 0.598736 0.73744 0.732328
k2v2 0.334262 0.525527 0.760021 0.72138 0.757424 0.766938 0.801972 0.808949
k2v4 0.333106 0.558613 0.765639 0.732337 0.762769 0.760775 0.799323 0.808594
k2v6 0.339966 0.564876 0.764105 0.730226 0.763797 0.765849 0.797442 0.80742
k2v8 0.331698 0.567622 0.763972 0.733084 0.765828 0.769881 0.79908 0.803599
k4v1 0.496626 0.575004 0.768665 0.803298 0.773274 0.815164 0.823299 0.817609
k4v2 0.666081 0.693266 0.797161 0.805846 0.785901 0.822081 0.819957 0.819761
k4v4 0.691803 0.701052 0.797872 0.806737 0.788327 0.818848 0.816583 0.824813
k4v6 0.688022 0.701648 0.797325 0.809305 0.784273 0.817312 0.818533 0.823784
k4v8 0.686547 0.700206 0.79702 0.807328 0.782431 0.817872 0.818202 0.822701
k6v1 0.649633 0.597133 0.770615 0.801156 0.765218 0.812864 0.817002 0.816527
k6v2 0.692575 0.699732 0.798369 0.80343 0.784922 0.820546 0.821854 0.825656
k6v4 0.701547 0.705158 0.79613 0.806657 0.786 0.820736 0.822193 0.824622
k6v6 0.700741 0.705549 0.795475 0.807525 0.790379 0.81943 0.821026 0.825455
k6v8 0.693932 0.702742 0.79642 0.806614 0.789155 0.81943 0.820462 0.825005
k8v1 0.650554 0.598797 0.768704 0.802301 0.767131 0.814654 0.818909 0.818461
k8v2 0.691099 0.696408 0.798332 0.802632 0.787855 0.822552 0.822365 0.824697
k8v4 0.70116 0.703779 0.794958 0.807328 0.788884 0.820734 0.822508 0.825976
k8v6 0.698706 0.70461 0.795362 0.808917 0.786832 0.818963 0.820002 0.825695
k8v8 0.696944 0.70377 0.794663 0.808457 0.785888 0.820653 0.819874 0.826971
bf16 0.707878 0.70117 0.794959 0.807871 0.785389 0.820364 0.81985 0.826088
Table 8: Full EQ-Bench Parseable.
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-32B
k1v1 0 0 0 0 0 0 0
k1v2 0 0 0 0 0 0 0
k1v4 0 0 0 0 0 0 0
k1v6 0 0 0 0 0 0 0
k1v8 0 0 0 0 0 0 0
k2v1 0 0 8.187 0 35.673 0 2.339
k2v2 0 16.374 80.702 27.485 85.965 65.497 67.836
k2v4 0 34.503 84.795 39.766 88.889 70.175 70.175
k2v6 0 39.181 83.041 36.842 90.059 67.836 69.591
k2v8 0 38.012 84.211 40.936 88.889 68.421 70.760
k4v1 0 47.953 58.480 77.778 83.626 98.830 77.778
k4v2 63.743 97.076 95.322 87.719 97.661 95.322 80.702
k4v4 65.497 97.076 98.830 84.795 99.415 95.322 80.702
k4v6 64.328 97.076 98.830 85.380 98.830 93.567 80.702
k4v8 62.573 97.076 98.830 84.795 98.830 93.567 80.702
k6v1 23.392 60.234 53.801 81.287 90.643 96.491 79.532
k6v2 99.415 97.076 94.737 87.719 97.661 94.737 80.702
k6v4 99.415 97.661 99.415 87.135 98.830 95.906 80.702
k6v6 99.415 97.661 98.830 86.550 98.830 96.491 80.702
k6v8 99.415 97.661 98.830 87.135 98.830 97.661 80.702
k8v1 26.901 60.234 56.725 78.947 92.983 97.076 77.193
k8v2 99.415 97.076 94.737 87.135 97.661 96.491 79.532
k8v4 99.415 97.661 99.415 87.719 98.830 95.322 80.702
k8v6 99.415 97.661 99.415 86.550 98.830 95.906 80.702
k8v8 99.415 97.661 99.415 86.550 98.830 95.906 80.702
bf16 100 97.661 99.415 87.719 98.830 94.737 79.532
Table 9: Full GSM8K Accuracies (Flexible).
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0.0106141 0.0121304 0.0159212 0.00985595 0.0159212 0.0106141 0.0144049 0.0204701
k1v2 0.00833965 0.0144049 0.015163 0.00985595 0.0197119 0.0128886 0.0174375 0.0181956
k1v4 0.0090978 0.0197119 0.025019 0.0144049 0.0204701 0.0075815 0.00985595 0.0106141
k1v6 0.00985595 0.0106141 0.0272934 0.0174375 0.0197119 0.0075815 0.0045489 0.0159212
k1v8 0.00833965 0.015163 0.0280516 0.0090978 0.015163 0.0136467 0.0113723 0.0106141
k2v1 0.0030326 0.0181956 0.015163 0.00833965 0.0500379 0.0128886 0.0204701 0.0144049
k2v2 0.0030326 0.0288097 0.278241 0.0265353 0.38514 0.128886 0.327521 0.38514
k2v4 0.0030326 0.0197119 0.351024 0.0432146 0.45489 0.170584 0.412434 0.43442
k2v6 0.00379075 0.0257771 0.3442 0.040182 0.445792 0.169067 0.396513 0.419257
k2v8 0.00379075 0.0280516 0.354814 0.0348749 0.460197 0.166793 0.394238 0.448825
k4v1 0.0227445 0.0727824 0.281274 0.576952 0.517058 0.77862 0.70887 0.721001
k4v2 0.040182 0.294162 0.639121 0.808188 0.756634 0.87718 0.804397 0.648976
k4v4 0.040182 0.332828 0.652767 0.830174 0.777862 0.87718 0.823351 0.629265
k4v6 0.055345 0.33207 0.66793 0.831691 0.762699 0.88097 0.828658 0.619409
k4v8 0.0416983 0.341168 0.658074 0.826384 0.75815 0.877938 0.835481 0.631539
k6v1 0.0932525 0.0962851 0.307809 0.632297 0.521607 0.795299 0.742987 0.701289
k6v2 0.367703 0.3116 0.644428 0.830174 0.761941 0.880212 0.824867 0.59818
k6v4 0.404852 0.333586 0.653525 0.836998 0.770281 0.877938 0.827142 0.610311
k6v6 0.420015 0.336619 0.662623 0.843821 0.771797 0.884003 0.8279 0.595906
k6v8 0.41395 0.341926 0.664898 0.840788 0.771039 0.886277 0.826384 0.608036
k8v1 0.0879454 0.0955269 0.304018 0.636846 0.529947 0.797574 0.742987 0.708112
k8v2 0.389689 0.313116 0.64746 0.825625 0.74981 0.882487 0.822593 0.605762
k8v4 0.416224 0.340409 0.660349 0.832449 0.770281 0.874905 0.831691 0.620925
k8v6 0.419257 0.341926 0.667172 0.843063 0.767248 0.880212 0.839272 0.608036
k8v8 0.413192 0.333586 0.666414 0.843063 0.758908 0.880212 0.839272 0.603487
bf16 0.412434 0.328279 0.663381 0.845337 0.759666 0.87718 0.835481 0.603487
Table 10: Full GSM8K Accuracies (Strict).
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0 0 0 0 0 0 0 0
k1v2 0 0 0 0 0 0 0 0
k1v4 0 0 0 0 0 0 0.00075815 0
k1v6 0 0 0.0015163 0 0 0.00075815 0 0
k1v8 0 0 0 0 0 0 0.00075815 0
k2v1 0 0.00075815 0.00530705 0 0.0280516 0 0.00075815 0.0015163
k2v2 0 0.0113723 0.275967 0.00379075 0.379833 0.0667172 0.260045 0.234268
k2v4 0 0.00985595 0.350265 0.0090978 0.443518 0.103867 0.343442 0.334344
k2v6 0 0.0121304 0.3442 0.00985595 0.438969 0.115239 0.348749 0.329795
k2v8 0 0.0136467 0.353298 0.00530705 0.448825 0.109174 0.35254 0.347233
k4v1 0.0060652 0.0614102 0.292646 0.64746 0.500379 0.785444 0.824867 0.701289
k4v2 0.0424564 0.297195 0.627748 0.824867 0.750569 0.875663 0.882487 0.726308
k4v4 0.0439727 0.331312 0.642911 0.843821 0.758908 0.875663 0.877938 0.733889
k4v6 0.055345 0.331312 0.656558 0.839272 0.747536 0.880212 0.887036 0.733131
k4v8 0.0485216 0.341926 0.648976 0.834723 0.745262 0.87718 0.887036 0.736922
k6v1 0.0833965 0.0826384 0.326005 0.698256 0.505686 0.802881 0.830174 0.702047
k6v2 0.369219 0.307809 0.639121 0.839272 0.74981 0.879454 0.884003 0.70887
k6v4 0.404852 0.333586 0.643669 0.844579 0.755876 0.876422 0.884761 0.718726
k6v6 0.421531 0.335102 0.6558 0.852919 0.751327 0.881729 0.884003 0.71721
k6v8 0.41395 0.341168 0.657316 0.84837 0.752843 0.884761 0.884003 0.721001
k8v1 0.0849128 0.0841547 0.32373 0.704321 0.512509 0.805155 0.829416 0.695224
k8v2 0.388173 0.313874 0.639879 0.829416 0.739196 0.881729 0.886277 0.71721
k8v4 0.416983 0.340409 0.652009 0.845337 0.755876 0.874905 0.88097 0.721001
k8v6 0.419257 0.339651 0.661107 0.84837 0.749052 0.878696 0.886277 0.722517
k8v8 0.415466 0.33207 0.659591 0.850644 0.743745 0.876422 0.886277 0.723275
bf16 0.413192 0.327521 0.657316 0.852161 0.741471 0.874147 0.890826 0.717968